OpenStudy (anonymous):

1.Mac's farm has a total of 18 horses and 27 cows. What is the ratio of horses to cows in simplest form? a) 6/9 b)18/27 c)2/3 d)3/2 2.A class has 27 students. Fifteen are boys and the rest are girls. What is the ratio of girls to total students? a)12:15 b)15:12 c)27:12 d)12:27

OpenStudy (anonymous):

@dmezzullo @Ashleyisakitty

OpenStudy (anonymous):

@ash2326 help plz!

OpenStudy (anonymous):

1-c 2-a

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

no i think iam wrong

OpenStudy (anonymous):

ok :p

OpenStudy (anonymous):

let me think real quick

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

18/27

OpenStudy (anonymous):

For the first one just Up to a Fraction with you number . Look at this examlpe : If : Cows = 8 horses = 2 cows/horses = 8/2 = 4 Now I think you can answer :) For the second one : 27-15 = girls And help from last one :)

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

ok thx :)

OpenStudy (anonymous):

b

OpenStudy (anonymous):

b

OpenStudy (anonymous):

both are bs

OpenStudy (anonymous):

so 1-b 2-b

OpenStudy (anonymous):

@xdmb ! For the second said total students ! Think more ;)

OpenStudy (anonymous):

oh wow thx man

OpenStudy (anonymous):

;) No prob :)